A body of mass m moving with velocity V collides head-on with another body of mass 2m which is initially at rest. The ratio of K.E of colliding body before and after collision will be:

Grade:Select Grade

1 Answers

noogler
8 years ago
3:1
as the particles undergom head on collision first and second body moves with cmmon velocity
 
mv+2m(0)=(m+2m)vl ….......................by the conservation of momentum
vl=v/3
 
initial K.E............1/2mv2
final K.E..............1/2(m+2m)(vl)2=1/2(3m)(v/3)2=1/6mv2
K.E/K.Ef=3/1